# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 2.1. The composition product
By Proposition 7.9 there is a right adjoint symmetric monoidal embedding ${\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ into a symmetric monoidal $\infty$-category compatible with small colimits that preserves the initial object. So we can form the monoidal $\infty$-category ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})^{\otimes}\to{\mathrm{Ass}}$ and define ${\mathrm{sSeq}}({\mathcal{C}})^{\otimes}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})^{\otimes}$ as the full non-symmetric suboperad spanned by ${\mathrm{sSeq}}({\mathcal{C}}).$ Since the embedding ${\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ admits a left adjoint ${\mathrm{L}}$, the embedding ${\mathrm{sSeq}}({\mathcal{C}})\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})$ admits a left adjoint ${\mathrm{L}}_{*}$. Thus by Lemma 7.11 the restriction ${\mathrm{sSeq}}({\mathcal{C}})^{\otimes}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})^{\otimes}\to{\mathrm{Ass}}$ is a locally cocartesian fibration, where the composition product of ${\mathcal{O}}_{1},...,{\mathcal{O}}_{\ell}$ is ${\mathrm{L}}_{*}({\mathcal{O}}_{1}\circ...\circ{\mathcal{O}}_{\ell})$. So by Remark 2.7 we obtain a canonical equivalence
${\mathrm{L}}_{*}({\mathcal{O}}_{1}\circ...\circ{\mathcal{O}}_{\ell})_{\mathrm{n}}\simeq{\mathrm{L}}(({\mathcal{O}}_{1}\circ...\circ{\mathcal{O}}_{\ell})_{\mathrm{n}})\simeq$
$\mathrm{colim}_{\mathsf{f}\in{\mathrm{{\mathcal{F}}in}}^{\ell}_{\mathrm{n}}}{\mathrm{L}}(({\mathcal{O}}_{1})_{{\mathrm{n}}_{1}}\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{n}}_{1}}({\mathcal{O}}_{2})_{\mathsf{f}_{1}^{-1}({\mathrm{i}})}\otimes...\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{n}}_{\ell-1}}({\mathcal{O}}_{\ell})_{\mathsf{f}_{\ell-1}^{-1}({\mathrm{i}})})$
$\simeq\mathrm{colim}_{\mathsf{f}\in{\mathrm{{\mathcal{F}}in}}^{\ell}_{\mathrm{n}}}(({\mathcal{O}}_{1})_{{\mathrm{n}}_{1}}\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{n}}_{1}}({\mathcal{O}}_{2})_{\mathsf{f}_{1}^{-1}({\mathrm{i}})}\otimes...\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{n}}_{\ell-1}}({\mathcal{O}}_{\ell})_{\mathsf{f}_{\ell-1}^{-1}({\mathrm{i}})}).$
@@
2.2. $\infty$-operads and $\infty$-cooperads
@@
${\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}\subset{\mathrm{Op}}({\mathcal{C}})$
@@
$\widetilde{{\mathrm{Op}}}({\mathcal{C}})\subset{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}$
@@
$\mathrm{coOp}({\mathcal{C}})^{\mathrm{nu}}\subset\mathrm{coOp}({\mathcal{C}})$
@@
$\widetilde{\mathrm{coOp}}({\mathcal{C}})\subset\mathrm{coOp}({\mathcal{C}})^{\mathrm{nu}}$
@@
${\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}}):=\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}}^{\mathrm{op}})^{\mathrm{op}}.$
@@
Let ${\mathcal{C}}$ be a symmetric monoidal $\infty$-category compatible with small colimits. Every $\infty$-operad in ${\mathcal{C}}$ has an associated monad on ${\mathcal{C}}:$ the left action of the composition product on ${\mathrm{sSeq}}({\mathcal{C}})$ on ${\mathcal{C}}$ is the pullback of the endomorphism left action of $\mathrm{Fun}({\mathcal{C}},{\mathcal{C}})$ on ${\mathcal{C}}$ along a unique monoidal functor ${\mathrm{sSeq}}({\mathcal{C}})\to\mathrm{Fun}({\mathcal{C}},{\mathcal{C}})$ sending an $\infty$-operad ${\mathcal{O}}$ in ${\mathcal{C}}$ to a monad ${\mathrm{T}}_{\mathcal{O}}={\mathcal{O}}\circ(-)$ on ${\mathcal{C}}.$ Hence $\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$ identifies with ${\mathrm{LMod}}_{{\mathrm{T}}_{\mathcal{O}}}({\mathcal{C}})$, the $\infty$-category of ${\mathrm{T}}_{\mathcal{O}}$-algebras. Thus the monad associated to the free-forgetful adjunction ${\mathcal{C}}\rightleftarrows\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$ identifies with the monad associated to the free-forgetful adjunction ${\mathcal{C}}\rightleftarrows{\mathrm{LMod}}_{{\mathrm{T}}_{\mathcal{O}}}({\mathcal{C}})$, which is ${\mathrm{T}}_{\mathcal{O}}$ by [23, Proposition 4.7.3.3. 2.].
@@
1.: Since the symmetric monoidal structure on ${\mathcal{C}}$ is compatible with small colimits, the composition product on ${\mathrm{sSeq}}({\mathcal{C}})$ defines a monoidal structure whose tensor unit is ${\mathrm{triv}}.$ Hence by [23, Proposition 3.2.1.8., Lemma 3.2.1.10.] the $\infty$-category ${\mathrm{Op}}({\mathcal{C}})$ admits an initial object lying over ${\mathrm{triv}}.$
2.: Let $\Sigma_{\geq 2}:=\coprod_{{\mathrm{n}}\geq 2}{\mathrm{B}}\Sigma_{\mathrm{n}}$ and ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 2}:=\mathrm{Fun}(\Sigma_{\geq 2},{\mathcal{C}})\simeq\prod_{{\mathrm{n}}\geq 2}\mathrm{Fun}({\mathrm{B}}\Sigma_{\mathrm{n}},{\mathcal{C}}).$ Let
@@
3. follows from 1. and 2. If ${\mathcal{C}}$ is compatible with small colimits, the composition product on ${\mathrm{sSeq}}({\mathcal{C}})$ defines a monoidal structure. So the forgetful functor $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\to{\mathcal{C}}$ admits a left adjoint sending ${\mathrm{X}}$ to ${\mathcal{O}}\circ{\mathrm{X}}\simeq\coprod_{{\mathrm{n}}\geq 0}({\mathcal{O}}_{\mathrm{n}}\otimes{\mathrm{X}}^{\otimes{\mathrm{n}}})_{\Sigma_{\mathrm{n}}}.$ Hence $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})$ admits an initial object that lies over ${\mathcal{O}}_{0}.$ Moreover if ${\mathcal{C}}$ has a final object, $\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})={\mathrm{LMod}}_{\mathcal{O}}({\mathcal{C}})$ as a final object preserved by the forgetful functor to ${\mathcal{C}}.$
@@
1. follows from [23, Corollary 4.2.3.3.] and 2. follows from [23, Corollary 4.2.3.5.] since $\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$ is the $\infty$-category of left ${\mathcal{O}}$-modules with respect to the left action of the monoidal $\infty$-category ${\mathrm{sSeq}}({\mathcal{C}})$ on ${\mathcal{C}}$, where the left action functor ${\mathcal{O}}\circ(-):{\mathcal{C}}\to{\mathcal{C}}$ preserves small sifted colimits.
@@
For ${\mathrm{X}},{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})$ the structure morphism $\iota({\mathrm{X}})\otimes\iota({\mathrm{Y}})\to\iota({\mathrm{X}}\otimes{\mathrm{Y}})$ in ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})$ of $\iota$ induces in degree ${\mathrm{n}}\in\Sigma$ the canonical morphism
$\alpha:\underset{{\mathrm{k}}\in{\mathbb{N}}}{\coprod}(\underset{\underset{{\mathrm{n}}_{\mathrm{i}}\cap{\mathrm{n}}_{\mathrm{j}}=\emptyset,{\mathrm{i}}\neq{\mathrm{j}}}{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}={\mathrm{n}}}}{\coprod}{\mathrm{X}}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}))_{\Sigma_{\mathrm{k}}}\to\underset{{\mathrm{k}}\in{\mathbb{N}}}{\prod}(\underset{\underset{{\mathrm{n}}_{\mathrm{i}}\cap{\mathrm{n}}_{\mathrm{j}}=\emptyset,{\mathrm{i}}\neq{\mathrm{j}}}{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}={\mathrm{n}}}}{\prod}{\mathrm{X}}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}))^{\Sigma_{\mathrm{k}}}$
@@
We show that for any ${\mathrm{k}},{\mathrm{n}}\in\Sigma$ and ${\mathrm{X}},{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})$ such that ${\mathrm{Y}}_{0}$ is initial both maps
$\phi:(\underset{\underset{{\mathrm{n}}_{\mathrm{i}}\cap{\mathrm{n}}_{\mathrm{j}}=\emptyset,{\mathrm{i}}\neq{\mathrm{j}}}{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}={\mathrm{n}}}}{\coprod}{\mathrm{X}}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}))_{\Sigma_{\mathrm{k}}}\xrightarrow{}(\underset{\underset{{\mathrm{n}}_{\mathrm{i}}\cap{\mathrm{n}}_{\mathrm{j}}=\emptyset,{\mathrm{i}}\neq{\mathrm{j}}}{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}={\mathrm{n}}}}{\coprod}{\mathrm{X}}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}))^{\Sigma_{\mathrm{k}}},$
$\phi^{\prime}:(\underset{\underset{{\mathrm{n}}_{\mathrm{i}}\cap{\mathrm{n}}_{\mathrm{j}}=\emptyset,{\mathrm{i}}\neq{\mathrm{j}}}{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}={\mathrm{n}}}}{\coprod}{\mathrm{X}}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}))^{\Sigma_{\mathrm{k}}}\xrightarrow{}(\underset{\underset{{\mathrm{n}}_{\mathrm{i}}\cap{\mathrm{n}}_{\mathrm{j}}=\emptyset,{\mathrm{i}}\neq{\mathrm{j}}}{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}={\mathrm{n}}}}{\prod}{\mathrm{X}}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}))^{\Sigma_{\mathrm{k}}}$
@@
Let ${\mathrm{W}}$ be the set of ${\mathrm{k}}$-tuples $({\mathrm{n}}_{1},...,{\mathrm{n}}_{\mathrm{k}})$ of finite pairwise disjoint sets such that ${\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}={\mathrm{n}}$. Let ${\mathrm{W}}^{\prime}\subset{\mathrm{W}}$ be the subset of ${\mathrm{k}}$-tuples $({\mathrm{n}}_{1},...,{\mathrm{n}}_{\mathrm{k}})$ such that ${\mathrm{n}}_{1},...,{\mathrm{n}}_{\mathrm{k}}$ are not empty. Since ${\mathrm{W}}^{\prime}$ is finite, $\phi^{\prime}$ is an equivalence by preadditivity. To prove that $\phi$ is an equivalence, it is enough to check that the canonical $\Sigma_{\mathrm{k}}$-action on $\underset{\underset{{\mathrm{n}}_{\mathrm{i}}\cap{\mathrm{n}}_{\mathrm{j}}=\emptyset,{\mathrm{i}}\neq{\mathrm{j}}}{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}={\mathrm{n}}}}{\coprod}\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}$ is free because the preadditivity of ${\mathcal{C}}$ guarantees that free and cofree $\Sigma_{\mathrm{k}}$-actions coincide.
The set ${\mathrm{W}}$ carries a canonical $\Sigma_{\mathrm{k}}$-action such that the canonical embedding ${\mathrm{W}}\subset\Sigma^{\times{\mathrm{k}}}$ is $\Sigma_{\mathrm{k}}$-equivariant, where $\Sigma^{\times{\mathrm{k}}}$ carries the permutation action. The $\Sigma_{\mathrm{k}}$-action on ${\mathrm{W}}$ restricts to a free $\Sigma_{\mathrm{k}}$-action on ${\mathrm{W}}^{\prime}$ since the $\Sigma_{\mathrm{k}}$-action on ${\mathrm{W}}^{\prime}$ does not have fixed points. Composing with the $\Sigma_{\mathrm{k}}$-equivariant functor $\Sigma^{\times{\mathrm{k}}}\xrightarrow{{\mathrm{Y}}^{\times{\mathrm{k}}}}{\mathcal{C}}^{\prime\times{\mathrm{k}}}\to{\mathcal{C}}^{\prime}$ the map ${\mathrm{W}}^{\prime}\to{\mathrm{W}}$ is $\Sigma_{\mathrm{k}}$-equivariant as functor over ${\mathcal{C}}^{\prime}$, where ${\mathrm{W}}^{\prime}$ is free as category over ${\mathcal{C}}^{\prime}$. Taking colimits we obtain the following $\Sigma_{\mathrm{k}}$-equivariant equivalence, where the left hand side is free:
$\underset{\underset{{\mathrm{n}}_{\mathrm{i}}\cap{\mathrm{n}}_{\mathrm{j}}=\emptyset,{\mathrm{i}}\neq{\mathrm{j}},{\mathrm{n}}_{\mathrm{i}}\neq\emptyset}{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}={\mathrm{n}}}}{\coprod}\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}\simeq\underset{\underset{{\mathrm{n}}_{\mathrm{i}}\cap{\mathrm{n}}_{\mathrm{j}}=\emptyset,{\mathrm{i}}\neq{\mathrm{j}}}{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}={\mathrm{n}}}}{\coprod}\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}.$
@@
${\mathrm{coAlg}}({\mathrm{sSeq}}({\mathcal{C}}))\simeq\mathrm{coOp}({\mathcal{C}})$
@@
${\mathrm{coAlg}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\simeq\mathrm{coOp}({\mathcal{C}})^{\mathrm{nu}}.$
@@
${\mathrm{coAlg}}({\mathrm{sSeq}}({\mathcal{C}}))\to{\mathrm{coOp}}({\mathcal{C}})=\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}}^{\mathrm{op}}))^{\mathrm{op}}$
@@
${\mathrm{coAlg}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\simeq\mathrm{CoOp}({\mathcal{C}})^{\mathrm{nu}}=\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}}^{\mathrm{op}})_{\geq 1})^{\mathrm{op}}.$

### v2

@@ 2.1. The composition product
By Proposition 5.9 there is a right adjoint symmetric monoidal embedding ${\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ into a symmetric monoidal $\infty$-category compatible with small colimits that preserves the initial object. So we can form the monoidal $\infty$-category ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})^{\otimes}\to{\mathrm{Ass}}$ and define ${\mathrm{sSeq}}({\mathcal{C}})^{\otimes}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})^{\otimes}$ as the full non-symmetric suboperad spanned by ${\mathrm{sSeq}}({\mathcal{C}}).$ Since the embedding ${\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ admits a left adjoint $\mathrm{L}$}, the embedding ${\mathrm{sSeq}}({\mathcal{C}})\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})$ admits a left adjoint $\mathrm{L}$$_{*}$. Thus by [29, Lemma 3.13.] the restriction ${\mathrm{sSeq}}({\mathcal{C}})^{\otimes}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})^{\otimes}\to{\mathrm{Ass}}$ is a locally cocartesian fibration, where the composition product of ${\mathcal{O}}_{1},...,{\mathcal{O}}_{\ell}$ is ${\mathrm{L}}_{*}({\mathcal{O}}_{1}\circ...\circ{\mathcal{O}}_{\ell})$. So by Remark 2.7 we obtain a canonical equivalence
${\mathrm{L}}_{*}({\mathcal{O}}_{1}\circ...\circ{\mathcal{O}}_{\ell})_{n}\simeq{\mathrm{L}}(({\mathcal{O}}_{1}\circ...\circ{\mathcal{O}}_{\ell})_{n})\simeq$
$\mathrm{colim}_{{\mathrm{f}}\in{\mathrm{Fin}}^{\ell}_{n}}{\mathrm{L}}(({\mathcal{O}}_{1})_{{n}_{1}}\otimes\bigotimes_{{i}\in{n}_{1}}({\mathcal{O}}_{2})_{{\mathrm{f}}_{1}^{-1}({i})}\otimes...\otimes\bigotimes_{{i}\in{n}_{\ell-1}}({\mathcal{O}}_{\ell})_{{\mathrm{f}}_{\ell-1}^{-1}({i})})$
$\simeq\mathrm{colim}_{{\mathrm{f}}\in{\mathrm{Fin}}^{\ell}_{n}}(({\mathcal{O}}_{1})_{{n}_{1}}\otimes\bigotimes_{{i}\in{n}_{1}}({\mathcal{O}}_{2})_{{\mathrm{f}}_{1}^{-1}({i})}\otimes...\otimes\bigotimes_{{i}\in{n}_{\ell-1}}({\mathcal{O}}_{\ell})_{{\mathrm{f}}_{\ell-1}^{-1}({i})}).$
@@
$2.$ $\infty$-operads and $\infty$-cooperads
@@
${\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}\subset{\mathrm{Op}}({\mathcal{C}})$
@@
$\widetilde{{\mathrm{Op}}}({\mathcal{C}})\subset{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}$
@@
$\mathrm{coOp}({\mathcal{C}})^{\mathrm{nun}}\subset\mathrm{coOp}({\mathcal{C}})$
@@
$\widetilde{\mathrm{coOp}}({\mathcal{C}})\subset\mathrm{coOp}({\mathcal{C}})^{\mathrm{nun}}$
@@
$\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}}):=\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}}^{\mathrm{op}})^{\mathrm{op}}.$
@@
Let $\mathcal{C}$} be a symmetric monoidal $\infty$-category compatible with small colimits. Every $\infty$-operad in $\mathcal{C}$} has an associated monad on $\mathcal{C}$}: the left action of the composition product on ${\mathrm{sSeq}}({\mathcal{C}})$ on $\mathcal{C}$} is the pullback of the endomorphism left action of $\mathrm{Fun}({\mathcal{C}},{\mathcal{C}})$ on $\mathcal{C}$} along a unique monoidal functor ${\mathrm{sSeq}}({\mathcal{C}})\to\mathrm{Fun}({\mathcal{C}},{\mathcal{C}})$ sending an $\infty$-operad $\mathcal{O}$} in $\mathcal{C}$} to a monad ${\mathrm{T}}_{\mathcal{O}}={\mathcal{O}}\circ(-)$ on $\mathcal{C}$}. Hence $\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$ identifies with ${\mathrm{LMod}}_{{\mathrm{T}}_{\mathcal{O}}}({\mathcal{C}})$, the $\infty$-category of ${\mathrm{T}}_{\mathcal{O}}$-algebras. Thus the monad associated to the free-forgetful adjunction ${\mathcal{C}}\rightleftarrows\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$ identifies with the monad associated to the free-forgetful adjunction ${\mathcal{C}}\rightleftarrows{\mathrm{LMod}}_{{\mathrm{T}}_{\mathcal{O}}}({\mathcal{C}})$, which is ${\mathrm{T}}_{\mathcal{O}}$ by [39, Proposition 4.7.3.3. 2.].
@@
1.: Since the symmetric monoidal structure on $\mathcal{C}$} is compatible with small colimits, the composition product on ${\mathrm{sSeq}}({\mathcal{C}})$ defines a monoidal structure whose tensor unit is $\mathrm{triv}$}. Hence by [39, Proposition 3.2.1.8., Lemma 3.2.1.10.] the $\infty$-category ${\mathrm{Op}}({\mathcal{C}})$ admits an initial object lying over $\mathrm{triv}$}.
2.: Let $\Sigma_{\geq 2}:=\coprod_{{n}\geq 2}{\mathrm{B}}\Sigma_{n}$ and ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 2}:=\mathrm{Fun}(\Sigma_{\geq 2},{\mathcal{C}})\simeq\prod_{{n}\geq 2}\mathrm{Fun}({\mathrm{B}}\Sigma_{n},{\mathcal{C}}).$ Let
@@
3. follows from 1. and 2. If $\mathcal{C}$} is compatible with small colimits, the composition product on ${\mathrm{sSeq}}({\mathcal{C}})$ defines a monoidal structure. So the forgetful functor $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\to{\mathcal{C}}$ admits a left adjoint sending $\mathrm{X}$} to ${\mathcal{O}}\circ{\mathrm{X}}\simeq\coprod_{{n}\geq 0}({\mathcal{O}}_{n}\otimes{\mathrm{X}}^{\otimes{n}})_{\Sigma_{n}}.$ Hence $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})$ admits an initial object that lies over ${\mathcal{O}}_{0}.$ Moreover if $\mathcal{C}$} has a final object, $\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})={\mathrm{LMod}}_{\mathcal{O}}({\mathcal{C}})$ as a final object preserved by the forgetful functor to $\mathcal{C}$}.
@@
1. follows from [39, Corollary 4.2.3.3.] and 2. follows from [39, Corollary 4.2.3.5.] since $\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$ is the $\infty$-category of left $\mathcal{O}$}-modules with respect to the left action of the monoidal $\infty$-category ${\mathrm{sSeq}}({\mathcal{C}})$ on $\mathcal{C}$}, where the left action functor ${\mathcal{O}}\circ(-):{\mathcal{C}}\to{\mathcal{C}}$ preserves small sifted colimits.
@@
For ${\mathrm{X}},{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})$ the structure morphism $\iota({\mathrm{X}})\otimes\iota({\mathrm{Y}})\to\iota({\mathrm{X}}\otimes{\mathrm{Y}})$ in ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})$ of $\iota$ induces in degree ${n}\in\Sigma$ the canonical morphism
$\alpha:\underset{{k}\in{\mathbb{N}}}{\coprod}(\underset{\underset{{n}_{i}\cap{n}_{j}=\emptyset,{i}\neq{j}}{{n}_{1}\coprod...\coprod{n}_{k}={n}}}{\coprod}{\mathrm{X}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}))_{\Sigma_{k}}\to\underset{{k}\in{\mathbb{N}}}{\prod}(\underset{\underset{{n}_{i}\cap{n}_{j}=\emptyset,{i}\neq{j}}{{n}_{1}\coprod...\coprod{n}_{k}={n}}}{\prod}{\mathrm{X}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}))^{\Sigma_{k}}$
@@
We show that for any ${k},{n}\in\Sigma$ and ${\mathrm{X}},{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})$ such that ${\mathrm{Y}}_{0}$ is initial both maps
$\phi:(\underset{\underset{{n}_{i}\cap{n}_{j}=\emptyset,{i}\neq{j}}{{n}_{1}\coprod...\coprod{n}_{k}={n}}}{\coprod}{\mathrm{X}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}))_{\Sigma_{k}}\xrightarrow{}(\underset{\underset{{n}_{i}\cap{n}_{j}=\emptyset,{i}\neq{j}}{{n}_{1}\coprod...\coprod{n}_{k}={n}}}{\coprod}{\mathrm{X}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}))^{\Sigma_{k}},$
$\phi^{\prime}:(\underset{\underset{{n}_{i}\cap{n}_{j}=\emptyset,{i}\neq{j}}{{n}_{1}\coprod...\coprod{n}_{k}={n}}}{\coprod}{\mathrm{X}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}))^{\Sigma_{k}}\xrightarrow{}(\underset{\underset{{n}_{i}\cap{n}_{j}=\emptyset,{i}\neq{j}}{{n}_{1}\coprod...\coprod{n}_{k}={n}}}{\prod}{\mathrm{X}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}))^{\Sigma_{k}}$
@@
Let $\mathrm{W}$} be the set of k{k}-tuples $({n}_{1},...,{n}_{k})$ of finite pairwise disjoint sets such that ${n}_{1}\coprod...\coprod{n}_{k}={n}$. Let ${\mathrm{W}}^{\prime}\subset{\mathrm{W}}$ be the subset of k{k}-tuples $({n}_{1},...,{n}_{k})$ such that ${n}_{1},...,{n}_{k}$ are not empty. Since $\mathrm{W}$$^{\prime}$ is finite, $\phi^{\prime}$ is an equivalence by preadditivity. To prove that $\phi$ is an equivalence, it is enough to check that the canonical $\Sigma_{k}$-action on $\underset{\underset{{n}_{i}\cap{n}_{j}=\emptyset,{i}\neq{j}}{{n}_{1}\coprod...\coprod{n}_{k}={n}}}{\coprod}\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}$ is free because the preadditivity of $\mathcal{C}$} guarantees that free and cofree $\Sigma_{k}$-actions coincide.
The set $\mathrm{W}$} carries a canonical $\Sigma_{k}$-action such that the canonical embedding ${\mathrm{W}}\subset\Sigma^{\times{k}}$ is $\Sigma_{k}$-equivariant, where $\Sigma^{\times{k}}$ carries the permutation action. The $\Sigma_{k}$-action on $\mathrm{W}$} restricts to a free $\Sigma_{k}$-action on $\mathrm{W}$$^{\prime}$ since the $\Sigma_{k}$-action on $\mathrm{W}$$^{\prime}$ does not have fixed points. Composing with the $\Sigma_{k}$-equivariant functor $\Sigma^{\times{k}}\xrightarrow{{\mathrm{Y}}^{\times{k}}}{\mathcal{C}}^{\prime\times{k}}\to{\mathcal{C}}^{\prime}$ the map ${\mathrm{W}}^{\prime}\to{\mathrm{W}}$ is $\Sigma_{k}$-equivariant as functor over $\mathcal{C}$$^{\prime}$, where $\mathrm{W}$$^{\prime}$ is free as category over $\mathcal{C}$$^{\prime}$. Taking colimits we obtain the following $\Sigma_{k}$-equivariant equivalence, where the left hand side is free:
$\underset{\underset{{n}_{i}\cap{n}_{j}=\emptyset,{i}\neq{j},{n}_{i}\neq\emptyset}{{n}_{1}\coprod...\coprod{n}_{k}={n}}}{\coprod}\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}\simeq\underset{\underset{{n}_{i}\cap{n}_{j}=\emptyset,{i}\neq{j}}{{n}_{1}\coprod...\coprod{n}_{k}={n}}}{\coprod}\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}.$
@@
$\mathrm{co}\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}}))\simeq\mathrm{coOp}({\mathcal{C}})$
@@
$\mathrm{co}\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\simeq\mathrm{coOp}({\mathcal{C}})^{\mathrm{nun}}.$
@@
$\mathrm{co}\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}}))\to{\mathrm{coOp}}({\mathcal{C}})=\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}}^{\mathrm{op}}))^{\mathrm{op}}$
@@
$\mathrm{co}\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\simeq\mathrm{CoOp}({\mathcal{C}})^{\mathrm{nun}}=\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}}^{\mathrm{op}})_{\geq 1})^{\mathrm{op}}.$